When dragons on planet Pern lay eggs, the eggs are either green or yellow. The biologists have observed over the years of the dragon eggs to incubate. Consider all the possible samples of 61 dragon eggs. of the eggs are yellow, and the rest green. Next spring the lead scientist has permission to randomly select 61 What is the asual number of yellow eggs in samples of 61 eggs? (Give answers as SENSIBLE whole numbers) minimum usual number of yellow eggs maximum usual number of yellow eggs =
Answer #1

Answer:

Mean = np = 61*(0.30) = 18.3

Standard deviation = sqrt(npq) = sqrt( 61 * 0.30 * 0.70) = 3.579

Hence,

Minimum usual value = 18.3 - 2*3.579 = 11.868 = 11.14

Maximum usual value = 18.3 + 2*3.579 = 26.532 = 25.458